OpenStudy (anonymous):

Find the exact value by using a half-angle identity.. tan 7pi/8

OpenStudy (anonymous):

\[ \tan \left(\frac{7 \pi }{8}\right)=\tan \left(\pi -\frac{\pi }{8}\right)=-\tan \left(\frac{\pi }{8}\right)=\\-\frac{\sin \left(\frac{\pi }{4}\right)}{1+\cos \left(\frac{\pi }{4}\right)}=-\frac{1}{\sqrt{2} \left(1+\frac{1}{\sqrt{2}}\right)}=1-\sqrt{2} \]

OpenStudy (anonymous):

We used the fact that \[ \tan\left( \frac u 2 \right)=\frac{\sin(u)}{1+\cos(u)} \]

OpenStudy (anonymous):

Thank you! I know what formula to use I just get stuck in the solving process once i substitute everything in...

OpenStudy (anonymous):

YW